# Variational Bayesian Bow tie Neural Networks with Shrinkage

### v2

@@ B.1 ELBO for training
Recall, that optimal variational parameters maximize the ELBO function of Equation 10 which for our model is:
ELBO $=\mathbb{E}\left[\log p(\mathbf{y},\mathbf{a},\bm{\gamma},\bm{\omega}|\mathbf{W},\mathbf{b},\bm{\Sigma})\right]+\mathbb{E}\left[\log p(\mathbf{W}|\bm{\psi},\bm{\tau})\right]+\mathbb{E}\left[\log p(\bm{\psi})\right]+\mathbb{E}\left[\log p(\bm{\tau})\right]$
$+\mathbb{E}\left[\log p(\mathbf{b})\right]+\mathbb{E}\left[\log p(\bm{\Sigma})\right]-\mathbb{E}\left[\log q(\mathbf{a})\right]-\mathbb{E}\left[\log q(\bm{\gamma})\right]-\mathbb{E}\left[\log q(\bm{\omega})\right]$
$-\mathbb{E}\left[\log q(\mathbf{W},\mathbf{b})\right]-\mathbb{E}\left[\log q(\bm{\eta})\right]-\mathbb{E}\left[\log q(\bm{\psi})\right]-\mathbb{E}\left[\log q(\bm{\tau})\right].$
@@
$=C_{\tau}+\sum_{l=1}^{L+1}\mathbb{E}\left[\log\tau_{l}^{\nu_{\text{glob}}-1}\exp\left(-\frac{1}{2}\left(\frac{\delta_{\text{glob}}^{2}}{\tau_{l}}+\lambda_{\text{glob}}^{2}\tau_{l}\right)\right)\right]-\mathbb{E}\left[\log\left(\tau_{l}^{\hat{\nu}_{\text{glob},l}-1}\right)\exp\left(-\frac{1}{2}\left(\frac{\hat{\delta}_{\text{glob},l}^{2}}{\tau_{l}}+\lambda_{\text{glob}}^{2}\tau_{l}\right)\right)\right]$
@@
$\mathbb{E}\left[\log p(\bm{\psi})-\log q(\bm{\psi})\right]=C_{\bm{\psi}}+\sum_{l=1}^{L+1}\sum_{d=1}^{D_{l}}\sum_{d^{\prime}=1}^{D_{l-1}}\mathbb{E}\left[\log\text{GIG}\left(\psi_{l,d,d^{\prime}}\mid\nu_{\text{loc},l},\delta_{\text{loc},l},\lambda_{\text{loc},l}\right)\right]$
$-\sum_{l=1}^{L+1}\sum_{d=1}^{D_{l}}\sum_{d^{\prime}=1}^{D_{l-1}}\mathbb{E}\left[\log\text{GIG}\left(\psi_{l,d,d^{\prime}}|\hat{\nu}_{\text{loc},l,d,d^{\prime}},\hat{\delta}_{\text{loc},l,d,d^{\prime}},\lambda_{\text{loc},l}\right)\right]$
$=C_{\bm{\psi}}+\sum_{l=1}^{L+1}\sum_{d=1}^{D_{l}}\sum_{d^{\prime}=1}^{D_{l-1}}\mathbb{E}\left[\log\psi_{l,d,d^{\prime}}^{\nu_{\text{loc},l}-1}\exp\left(-\frac{1}{2}\left(\frac{\delta_{\text{loc},l}^{2}}{\psi_{l,d,d^{\prime}}}+\lambda_{\text{loc},l}^{2}\psi_{l,d,d^{\prime}}\right)\right)\right]$
$-\sum_{l=1}^{L+1}\sum_{d=1}^{D_{l}}\sum_{d^{\prime}=1}^{D_{l-1}}\mathbb{E}\left[\log\left(\psi_{l,d,d^{\prime}}^{\hat{\nu}_{\text{loc},l,d,d^{\prime}}-1}\right)\exp\left(-\frac{1}{2}\left(\frac{\hat{\delta}_{\text{loc},l,d,d^{\prime}}^{2}}{\psi_{l,d,d^{\prime}}}+\lambda_{\text{loc},l}^{2}\psi_{l,d,d^{\prime}}\right)\right)\right]$
@@
As before, the covariance matrix matrix is assumed to be diagonal so that the relevant ELBO is:
@@
$=\sum_{n=1}^{N}\sum_{d=1}^{D_{L+1}}\mathbb{E}\left[\log\text{N}\left(y_{n,d}\mid\mathbf{z}_{n,L+1,d},\bm{\Sigma}_{L+1,d}\right)\right]+\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\log\text{N}\left(a_{n,l,d}\mid\bm{\gamma}_{n,d}\odot\mathbf{z}_{n,l,d},\bm{\Sigma}_{l,d}\right)\right]$
@@
$-\sum_{n=1}^{N}\sum_{l=1}^{L}\mathbb{E}\left[\log\text{N}\left(\mathbf{a}_{n,l}\mid\mathbf{t}_{n,l}+\mathbf{M}_{n,l}\mathbf{a}_{n,l-1},\mathbf{S}_{n,l}\right)\right]$
$-\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\log\text{Bern}\left(\gamma_{n,l,d}\mid\rho_{n,l,d}\right)\right]+\mathbb{E}\left[\log\text{PG}(\omega_{n,l,d}\mid 1,A_{n,l,d})\right]$
@@
$+\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\log(\eta_{l,d}^{2})^{-1/2}\exp\left(-\frac{1}{2\eta_{l,d}^{2}}\left(a_{n,l,d}-\gamma_{n,l,d}\odot\left(\mathbf{W}_{l,d}\mathbf{a}_{n,l-1}+b_{l,d}\right)\right)^{2}\right)\right]-N\sum_{l=1}^{L}D_{l}\log(2)$
@@
$-\frac{1}{2}\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\frac{1}{\eta_{l,d}^{2}}\right]\left(\mathbb{E}\left[a_{n,l,d}^{2}\right]-2\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}a_{n,l,d}\right]\right)$
@@
Then, we can sum the derived above parts to get the total ELBO of our model:
@@ B.2 ELBO for prediction
$=\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\log\text{N}\left(a_{*,l,d}\mid\bm{\gamma}_{*,d}\odot\mathbf{z}_{*,l,d},\bm{\Sigma}_{l,d}\right)\right]+\mathbb{E}\left[\log\left(\exp\left(\frac{\kappa_{*,l,d}z_{*,l,d}}{T}\right)\exp\left(-\frac{\omega_{*,l,d}z_{*,l,d}^{2}}{2T^{2}}\right)\text{PG}(\omega_{n,l,d}\mid 1,0)\right)\right]$

### v3

@@ B.1 ELBO for training
Recall that optimal variational parameters maximize the ELBO function of Equation 10, which for our model is:
$\text{ELBO}=\mathbb{E}\left[\log p(\mathbf{y},\mathbf{a},\bm{\gamma},\bm{\omega}|\mathbf{W},\mathbf{b},\bm{\Sigma})\right]+\mathbb{E}\left[\log p(\mathbf{W}|\bm{\psi},\bm{\tau})\right]+\mathbb{E}\left[\log p(\bm{\psi})\right]+\mathbb{E}\left[\log p(\bm{\tau})\right]+\mathbb{E}\left[\log p(\mathbf{b})\right]+\mathbb{E}\left[\log p(\bm{\Sigma})\right]$
$-\mathbb{E}\left[\log q(\mathbf{a})\right]-\mathbb{E}\left[\log q(\bm{\gamma})\right]-\mathbb{E}\left[\log q(\bm{\omega})\right]-\mathbb{E}\left[\log q(\mathbf{W},\mathbf{b})\right]-\mathbb{E}\left[\log q(\bm{\eta})\right]-\mathbb{E}\left[\log q(\bm{\psi})\right]-\mathbb{E}\left[\log q(\bm{\tau})\right].$
@@
$=C_{\tau}+\sum_{l=1}^{L+1}\mathbb{E}\left[\log\tau_{l}^{\nu_{\text{glob}}-1}\exp\left(-\frac{1}{2}\left(\frac{\delta_{\text{glob}}^{2}}{\tau_{l}}+\lambda_{\text{glob}}^{2}\tau_{l}\right)\right)\right]-\sum_{l=1}^{L+1}\mathbb{E}\left[\log\left(\tau_{l}^{\hat{\nu}_{\text{glob},l}-1}\right)\exp\left(-\frac{1}{2}\left(\frac{\hat{\delta}_{\text{glob},l}^{2}}{\tau_{l}}+\lambda_{\text{glob}}^{2}\tau_{l}\right)\right)\right]$
@@
$\mathbb{E}\left[\log p(\bm{\psi})-\log q(\bm{\psi})\right]=C_{\bm{\psi}}+\sum_{l=1}^{L+1}\sum_{d=1}^{D_{l}}\sum_{d^{\prime}=1}^{D_{l-1}}\mathbb{E}\left[\log\text{GIG}\left(\psi_{l,d,d^{\prime}}\mid\nu_{\text{loc},l},\delta_{\text{loc},l},\lambda_{\text{loc},l}\right)-\log\text{GIG}\left(\psi_{l,d,d^{\prime}}|\hat{\nu}_{\text{loc},l,d,d^{\prime}},\hat{\delta}_{\text{loc},l,d,d^{\prime}},\lambda_{\text{loc},l}\right)\right]$
$=C_{\bm{\psi}}+\sum_{l=1}^{L+1}\sum_{d=1}^{D_{l}}\sum_{d^{\prime}=1}^{D_{l-1}}\mathbb{E}\left[\log\psi_{l,d,d^{\prime}}^{\nu_{\text{loc},l}-1}\exp\left(-\frac{1}{2}\left(\frac{\delta_{\text{loc},l}^{2}}{\psi_{l,d,d^{\prime}}}+\lambda_{\text{loc},l}^{2}\psi_{l,d,d^{\prime}}\right)\right)-\log\left(\psi_{l,d,d^{\prime}}^{\hat{\nu}_{\text{loc},l,d,d^{\prime}}-1}\right)\exp\left(-\frac{1}{2}\left(\frac{\hat{\delta}_{\text{loc},l,d,d^{\prime}}^{2}}{\psi_{l,d,d^{\prime}}}+\lambda_{\text{loc},l}^{2}\psi_{l,d,d^{\prime}}\right)\right)\right]$
@@
As before, the covariance matrix is assumed to be diagonal so that the relevant ELBO is:
@@
$=\sum_{n=1}^{N}\sum_{d=1}^{D_{L+1}}\mathbb{E}\left[\log\text{N}\left(y_{n,d}\mid\mathbf{z}_{n,L+1,d},\bm{\Sigma}_{L+1,d}\right)\right]+\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\log\text{N}\left({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\mid\bm{\gamma}_{n,d}\odot\mathbf{z}_{n,l,d},\bm{\Sigma}_{l,d}\right)\right]$
@@
$-\sum_{n=1}^{N}\sum_{l=1}^{L}\mathbb{E}\left[\log\text{N}\left(\mathbf{a}_{n,l}\mid\mathbf{t}_{n,l}+\mathbf{M}_{n,l}\mathbf{a}_{n,l-1},\mathbf{S}_{n,l}\right)\right]-\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\log\text{Bern}\left(\gamma_{n,l,d}\mid\rho_{n,l,d}\right)\right]+\mathbb{E}\left[\log\text{PG}(\omega_{n,l,d}\mid 1,A_{n,l,d})\right]$
@@
$+\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\log(\eta_{l,d}^{2})^{-1/2}\exp\left(-\frac{1}{2\eta_{l,d}^{2}}\left({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}-\gamma_{n,l,d}\odot\left(\mathbf{W}_{l,d}\mathbf{a}_{n,l-1}+b_{l,d}\right)\right)^{2}\right)\right]-N\sum_{l=1}^{L}D_{l}\log(2)$
@@
$-\frac{1}{2}\sum_{n=1}^{N}\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\frac{1}{\eta_{l,d}^{2}}\right]\left(\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}^{2}\right]-2\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\right]\right)$
@@
Then, we can sum the derived parts above to get the total ELBO of our model:
@@ B.2 ELBO for prediction
$=\sum_{l=1}^{L}\sum_{d=1}^{D_{l}}\mathbb{E}\left[\log\text{N}\left(a_{*,l,d}\mid\bm{\gamma}_{*,d}\odot\mathbf{z}_{*,l,d},\bm{\Sigma}_{l,d}\right)\right]+\mathbb{E}\left[\log\left(\exp\left(\frac{\kappa_{*,l,d}z_{*,l,d}}{T}\right)\exp\left(-\frac{\omega_{*,l,d}z_{*,l,d}^{2}}{2T^{2}}\right)\text{PG}(\omega_{n,l,d}\mid 1,0)\right)\right]-$